OpenStudy (anonymous):

Prove that (x-a) is a factor of x^3 - (a + b + c)x^2 + (ab +bc + ca)x-abc. How do I prove this???

OpenStudy (anonymous):

show that if you replace \(x\) by \(a\) you get 0

OpenStudy (anonymous):

then you know if \(a\) is a root of a polynomial \(p(x)\) then you can factor as \(p(x)=(x-a)(q(x))\)

OpenStudy (anonymous):

yeah i tried that, but i couldnt end up with a zero..

OpenStudy (anonymous):

if you do not get 0, then you cannot factor

OpenStudy (anonymous):

i probably made a mistake.. because it has to equal to zero

OpenStudy (anonymous):

\[a^3 - (a + b + c)a^2 + (ab +bc + ca)a-abc\] is a startr

OpenStudy (anonymous):

yeah it works, try again

OpenStudy (anonymous):

a^3 - (a + b +c)a^2 + (ab+ bc +ca)a -abc = a^3 - a^3 + a^3b + a^3 c + a^2b + bc+ca^ - abc

OpenStudy (anonymous):

ohhh i see it, i messed up

OpenStudy (anonymous):

hmmm forgot the distributive property for the second term

OpenStudy (anonymous):

and some other mistakes too, but you can clean it up i am sure

OpenStudy (anonymous):

okay i tried and i ended up like this: 2a^2 + 2a^2c

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

ohh its negative.. i got it

OpenStudy (anonymous):

first step \[a^3 - (a + b + c)a^2 + (ab +bc + ca)a-abc\] second step ok

OpenStudy (anonymous):

yeah i didnt see the negative sign infront of the first bracket, thank you so much!!

OpenStudy (anonymous):

yw (you did all the work)